(i+under root 3)whole to the power100 +(i -under root 3)whole to the power 100 + 2 whole to the power 100=????

Grade:

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Gurpreet,

 

w = (-1 + sqrt.3 i)/2   or w = (-1 - sqrt.3 i)/2

i2 = -1

so, w = i (i + sqrt.3 )/2 and w = i (i - sqrt.3 )/2

put value of w = (i + sqrt.3 ) and w = (i - sqrt.3 ) in given quest. and use w3 = 1 and i2 = -1 to solve the problem.

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay